# Big Reasoning with Small Models: Instruction Retrieval at Inference Time

### v1

@@ Abstract
Can we bring big-scale reasoning to local-scale compute? Small language models (SLMs) are increasingly attractive for deployment because they run efficiently on local hardware, offering strong privacy, reduced cost, and lower environmental impact. Yet these models consistently struggle on tasks that require multi-step reasoning or specialized domain knowledge. To address this gap, we introduce instruction intervention at inference time. Our method builds an Instruction Corpus by grouping similar training questions and creating instructions that pair background knowledge with step-by-step procedures. At inference, the model retrieves the most relevant instructions and follows the steps. Unlike standard retrieval-augmented generation, which supplies raw passages, our approach provides structured scaffolds that guide reasoning. We evaluate this framework on MedQA (medical board exams), MMLU Professional Law, and MathQA using models from 3B to 14B parameters without any additional fine-tuning. Instruction retrieval yields consistent gains: 9.4% on MedQA, 7.9% on MMLU Law, and 5.1% on MathQA. Concise instructions outperform longer ones, and the magnitude of improvement depends strongly on model family and intrinsic reasoning ability.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 have demonstrated remarkable generalization and reasoning ability across a wide range of domains, from mathematical problem solving (Cobbe et al., 2021) and clinical diagnostics (Kwon et al., 2024) to legal and commonsense reasoning (Hendrycks et al., 2021a). Their success stems from massive parameter scale and diverse pretraining data, which allow them to internalize both factual knowledge and multi-step reasoning procedures. However, scaling to hundreds of billions of parameters introduces substantial costs. For example, serving a single 175B-parameter model requires over 300GB of GPU memory and specialized infrastructure (Frantar et al., 2023; Zheng et al., 2022), making real-time or on-prem deployment difficult. Moreover, production-grade LLMs typically operate as closed-source APIs, raising privacy and governance concerns in sensitive domains such as medicine and law. Small language models offer a practical alternative. SLMs can be deployed locally, audited directly, and fine-tuned for specific domains (Pham et al., 2024; Belcak et al., 2025; Fu et al., 2023). Their compact size lowers cost and latency while enabling operation on commodity hardware, including laptops and edge devices. Moreover, recent work (Srivastava et al., 2025) challenges the notion that reasoning is exclusive to large-scale models and that smaller models can achieve competitive reasoning performance when trained or compressed effectively. Yet, their reasoning behavior remains underexplored Zhu et al. (2024), and unlike LLMs, which encode broad world knowledge and reasoning patterns within their parameters, SLMs still have a limited capacity to internalize such information Fu et al. (2023). This gap motivates the central question of this work: how can small, efficient models achieve strong reasoning performance without scaling up their parameters or relying on additional training?
@@
Figure 1: Comparison of reasoning steps across model configurations on a MedQA-style case. The small language model alone produces an incorrect causal chain, while a large model reaches the right diagnosis but at high computational and privacy cost. Adding retrieved instructions supplies the missing background and procedural steps, enabling the SLM to reproduce the correct reasoning efficiently and privately.
@@
We address this question through instruction intervention at inference time, which augments small models with explicit, retrievable reasoning procedures. Instead of generating reasoning steps it cannot reliably produce on its own, the model retrieves structured instructions that pair domain background with step-by-step guidance. As illustrated in Figure 1, this approach externalizes reasoning as a retrieval process, supplying the scaffolds that SLMs lack while retaining their efficiency and privacy advantages. The method is entirely inference-based: a single, unmodified SLM retrieves and follows instructions without additional training, and the approach generalizes across domains.
@@
We evaluate this framework across three reasoning benchmarks (MedQA, MMLU Law, and MathQA) to assess both its effectiveness and the conditions under which it provides the greatest benefit. Specifically, we investigate two questions. First, does instruction retrieval reliably improve reasoning for small models, and what level of model capacity is needed for those gains to appear. Second, which properties of the retrieved instructions matter most. We test structure, length, and level of detail. We show that instruction retrieval improves accuracy by 5–10 percentage points for models above 3B parameters without additional fine-tuning. Concise, well-structured instructions that integrate factual knowledge with step-by-step reasoning yield the largest gains, and performance remains stable across clustering granularities, showing that the approach is robust to corpus design. Mixed-effects analysis indicates that improvement depends more on a model’s ability to follow structured guidance than on size. Notably, on knowledge-intensive tasks such as MedQA and MMLU Law, a 14B-parameter model equipped with retrieved instructions surpasses GPT-4o in zero-shot accuracy. Our results suggest that externalizing reasoning as retrievable text offers a practical path to scaling reliable inference on resource-limited or privacy-sensitive hardware. All code and the Instruction Corpus will be made publicly available upon acceptance to support reproducibility.
@@ 2 Related Work
Improving the reasoning ability of small language models has been approached from multiple directions. Chain-of-Thought (CoT) prompting encourages models to decompose a reasoning task into a sequence of intermediate steps rather than attempting to produce an answer directly. This approach has been shown to substantially improve the performance of LLMs across commonsense, symbolic, and mathematical reasoning benchmarks (Wei et al., 2023; Kojima et al., 2023; Wang et al., 2023). These gains, however, rely heavily on scale: only models with tens or hundreds of billions of parameters such as PaLM 540B (Chowdhery et al., 2022) or GPT-3 175B (Brown et al., 2020) reliably benefit from CoT prompting. Smaller models frequently generate illogical reasoning traces, and their accuracy can even decline when forced to produce step-by-step rationales (Wei et al., 2023).
@@
High School Concise 453 4.31 4.91 …See the big clue (underweight athlete with missed periods), match it to the rule (FHA $\rightarrow$ low estrogen), then pick “low bone density”…
Graduate Concise 631 4.57 4.92 …Prioritize the diagnostic triad (amenorrhea + low BMI + training), apply HPO suppression logic, reject violated conditions (e.g., high TSH/estrogen), and select “decreased bone density”…
@@
Graduate Verbose 1,765 4.97 4.69 …Synthesize the amenorrhea–energy-deficit pattern, map it via leptin/kisspeptin $\rightarrow$ GnRH suppression to hypoestrogenism, weigh alternative axes (thyroid, prolactin, PCOS) by contradiction, and justify “decreased bone density” as the compelled outcome…
@@
Table 1: Comparison of four instruction variants differing in audience level and length. Columns report the average token length and GPT-based quality scores (five-point Likert) for knowledge comprehensiveness and relevance, with example excerpts from MedQA. Concise variants are 2–3× shorter while retaining high relevance, whereas verbose variants expand coverage at the cost of slightly lower relevance.
@@
To overcome the scale dependence of prompting, a parallel line of work distills reasoning traces from LLMs into SLMs. These methods fine-tune smaller models on CoT-style rationales generated by larger models or derived from labeled data, with the goal of internalizing step-by-step reasoning skills (Ho et al., 2023; Li et al., 2022; Magister et al., 2023; Fu et al., 2023; Hsieh et al., 2023). Distillation has proven effective on arithmetic and symbolic reasoning benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), where reasoning patterns can be learned from repeated structures. However, the outcomes are less promising for knowledge-intensive tasks where accurate rationales depend on factual knowledge that smaller models struggle to retain. Because SLMs have limited parameter capacity, attempts to train them with extensive domain knowledge often lead to overspecialization and reduced generality (Fu et al., 2023).
@@
Rather than encoding all knowledge directly into model parameters, retrieval-based methods expand a model’s capacity by supplementing it with external information at inference time. Classical dense retrieval (Karpukhin et al., 2020) improves open-domain question answering by retrieving relevant passages, but the returned evidence is often noisy or unstructured, leaving the model to extract and organize reasoning steps on its own. More recent generate-then-retrieve frameworks address this limitation by synthesizing targeted context before retrieval, yielding cleaner and more relevant evidence for reasoning (Yu et al., 2023). In high-stakes domains such as medicine, curated knowledge bases can further improve reliability: retrieving from medical textbooks, for example, outperforms relying on generic sources like Wikipedia for clinical decision-making (Wang et al., 2025). While these approaches have been developed for LLMs, their lessons are even more pertinent for SLMs. Smaller models are more easily overloaded by long or noisy contexts and are less capable of abstracting structure from raw passages. As a result, the effectiveness of retrieval for SLMs depends on ensuring that the retrieved evidence is relevant and structured.
@@
Recent work (Kang et al., 2023; Zhao et al., 2024) improves upon unstructured retrieval by explicitly coupling evidence with the reasoning process itself. Knowledge-Augmented Reasoning Distillation (Kang et al., 2023) grounds LLM-style rationales in retrieved passages and employs a reranker to prioritize rationale-supporting evidence. This enables smaller models to learn reasoning tied to external knowledge, but still requires task-specific fine-tuning and keeps the reasoning policy embedded in parameters. Probe then Retrieve and Reason (Zhao et al., 2024) takes this modularity further by separating the pipeline into two distilled SLMs: a probing model that identifies what knowledge is needed and formulates retrieval queries, and a reasoning model that constructs step-by-step rationales from the retrieved passages. Our approach builds on the same intuition that retrieval should be directly tied to reasoning, but differs in where the reasoning resides. Rather than embedding procedures into model parameters or asking models to transform raw evidence into structured steps, we externalize reasoning into a reusable corpus of instructions. Our method largely separates reasoning from the model altogether.
@@ 3 Instruction Corpus
Small language models often fail on complex inference because they lack both sufficient domain knowledge and the structured support needed for multi-step reasoning. To address this gap, we construct the Instruction Corpus, a collection of modular instructions. Each instruction has two parts: (i) background knowledge relevant to a problem type and (ii) step-by-step reasoning procedures for solving it. At inference, instructions are retrieved and included in the prompt, guiding reasoning without task-specific fine-tuning. For example, in MedQA, a group of training questions may ask about contraindications for anticoagulants. From this group, we derive an instruction that outlines how to check mechanisms of action, review contraindications, and compare relative risks.
3.1 Corpus Construction
We construct the Instruction Corpus in three stages: clustering, instruction generation, and retrieval. First, training examples from each benchmark are embedded with OpenAI’s text-embedding-3-large and grouped using agglomerative clustering with average linkage and cosine distance. The resulting dendrogram supports different levels of granularity, where lower thresholds yield broad categories and higher thresholds produce highly specific clusters; by default, we adopt the most fine-grained partition so that each cluster corresponds to a distinct reasoning skill. Next, for each cluster we generate a reusable instruction by prompting GPT-5 with standardized templates (D) and up to 5 question examples from the cluster. Each instruction contains two sections: background knowledge summarizing the relevant domain concepts and a step-by-step procedure for solving new problems. The input examples serve only to guide generation and do not appear verbatim in the final instructions. Finally, at inference time, test queries are embedded with the same encoder and matched by cosine similarity to the closest clusters; the top-$k$ instructions (default $k=5$) are retrieved and included in prompt to provide both factual grounding and procedural scaffolding. See Appendix 2 for threshold values and statistics.
@@ 3.2 Instruction Variants
While instructions provide missing knowledge and reasoning scaffolds, their effectiveness depends on how they are written. We therefore study how variation in audience level and length shapes small-model performance. For the audience level, we create two variants. A graduate-level version preserves domain terminology and assumes advanced conceptual familiarity (e.g., clinical reasoning in MedQA). The high-school-level version simplifies vocabulary and logical complexity (e.g. explain this to me like a high schooler). Aligning readability with prior knowledge reduces intrinsic cognitive load and improves comprehension, a core prediction of Cognitive Load Theory (Renkl & Atkinson, 2003; Sweller, 2024). Along the depth axis, we contrast concise instructions that contain only the essential reasoning steps with longer instructions that include additional commentary to stimulate germane load and schema formation. Moreover, recent prompt-engineering studies show that prompt compression often preserves, and sometimes even enhances, downstream accuracy despite large token reductions, which underscores the practical value of concise variants (Renze & Guven, 2024; Li et al., 2024). Crossing the two factors yields four prompt variants per task. Descriptive quality summaries by style are reported in Appendix Table 4. We also include a baseline prompt that does not enforce any constraints on style or length. This version reflects the instructions produced directly from our generation pipeline and serves as a neutral reference point. Comparing the four controlled variants against this baseline help us to isolate the effects of readability and verbosity from artifacts of the generation process. Full prompt templates for all conditions, including the baseline, are provided in Appendix D.
3.3 Quality Validation
@@
Not all instructions are equally useful. Generic prompts such as “think carefully” may occasionally nudge performance but lack the structured knowledge and procedures that small models require. By contrast, the documents in our corpus are designed to provide both domain background and explicit reasoning steps. We verify that instructions are high quality and that the four variants differ only in style and length. To do so, we evaluate each instruction on three axes: knowledge (coverage and relevance of background facts), reasoning (soundness and task-specificity of steps), and clarity (structured, unambiguous presentation). A detailed rubric is given in Appendix A.1.3. Each document is scored automatically by GPT-5 on a five-point Likert scale. We repeat evaluations three times and report averaged scores; full results are shown in Table 4. Across tasks, mean scores exceed 4.7 on all three axes, confirming that the corpus provides accurate and reliable scaffolds rather than noisy or generic text. Table  1 summarizes quality differences across variants: verbosity increases comprehensiveness but slightly reduces relevance, while audience level has only minor effects. Clarity remains uniformly high across all conditions. Thus, the variants are equivalent in structure and quality, differing only in the controlled stylistic factors of length and audience.
3.4 Corpus Profile
@@
The Instruction Corpus varies systematically in both length and size across tasks. Concise instructions average 500–750 tokens, while verbose instructions expand to 1.3k–2.1k tokens, a two- to threefold increase in prompt length. Length distributions are shown in Appendix Figure 7, with per-task breakdowns in Figure 8. Graduate-level instructions are slightly longer and more detailed than high-school versions, but the influence of audience framing is minor compared to verbosity. At the default, fine-grained clustering threshold, the corpus contains 29k groups for MathQA, 8.4k for MedQA, and 1.0k for MMLU Law. Group sizes are highly skewed: often reducing to single examples, while the largest contain up to 12. Each produces a narrowly scoped instruction often grounded in only one or a few examples, highlighting the specific cues and reasoning steps needed for that problem type. Broader thresholds merge these singletons into larger groups, yielding more general instructions. We return to this question in Section 6, where we show that instruction retrieval remains robust across a wide range of cluster sizes. Full statistics and length distributions are reported in Appendix 2 and Appendix A.1.1.

### v2

@@ Abstract
Small language models (SLMs) enable low-cost, private, on-device inference, but they often fail on problems that require specialized domain knowledge or multi-step reasoning. Existing approaches for improving reasoning either rely on scale (e.g., chain-of-thought prompting), require task-specific training that limits reuse and generality (e.g., distillation), or retrieve unstructured information that still leaves the SLM to determine an appropriate reasoning strategy. We propose instruction retrieval, an inference-time intervention that augments an SLM with structured, reusable reasoning procedures rather than raw passages. We construct an Instruction Corpus by clustering similar training questions and using a teacher model to generate generalizable guides that pair domain background with explicit step-by-step procedures. At inference, the SLM retrieves the instructions most relevant to a given query and executes the associated procedures without any additional fine-tuning. Across three challenging domains—medicine, law, and mathematics, instruction retrieval yields consistent gains for models with at least 3B parameters, improving accuracy by 9.4%, 7.9%, and 5.1%, respectively, with the strongest 14B model surpassing GPT-4o’s zero-shot performance on knowledge-intensive tasks.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable generalization and reasoning ability across a wide range of domains, from mathematical problem solving (Cobbe et al., 2021) and clinical diagnostics (Kwon et al., 2024) to legal and commonsense reasoning (Hendrycks et al., 2021a). Much of their success stems from massive parameter scale and diverse pretraining data, which allow them to internalize both factual knowledge and multi-step reasoning procedures. However, scaling to hundreds of billions of parameters introduces substantial costs. For example, serving a single 175B-parameter model requires over 300GB of GPU memory and specialized infrastructure (Frantar et al., 2023; Zheng et al., 2022), making real-time or on-prem deployment difficult. Moreover, production-grade LLMs typically operate as closed-source APIs, raising privacy and governance concerns in sensitive domains such as medicine and law. Small language models offer a practical alternative. SLMs can be deployed locally, audited directly, and fine-tuned for specific domains (Pham et al., 2024; Belcak et al., 2025; Fu et al., 2023). Their compact size lowers cost and latency while enabling operation on commodity hardware, including laptops and edge devices. Moreover, recent work (Srivastava et al., 2025) challenges the notion that reasoning is exclusive to large-scale models and that smaller models can achieve competitive reasoning performance when trained or compressed effectively. Yet, their reasoning behavior remains underexplored Zhu et al. (2024), and unlike LLMs, which encode broad world knowledge and reasoning patterns within their parameters, SLMs still have a limited capacity to internalize such information Fu et al. (2023). This gap motivates the central question of this work: how can small, efficient models achieve strong reasoning performance without scaling up their parameters or relying on additional training for every new task?
@@
Figure 1: Comparison of reasoning steps across model configurations on a MedQA-style case. The small language model alone produces an incorrect chain, while a large model reaches the right diagnosis but at high computational and privacy cost. Adding retrieved instructions supplies the missing background and procedural steps, enabling the SLM to reproduce the correct reasoning efficiently and privately.
@@
We address this question through instruction intervention at inference time, which augments small models with explicit, retrievable reasoning procedures. Instead of generating reasoning steps it cannot reliably produce on its own, the model retrieves structured instructions generated by a larger model that pairs domain background with step-by-step guidance. As illustrated in Figure 1, this approach externalizes reasoning as a retrieval process, supplying the scaffolds that SLMs lack while retaining their efficiency and privacy advantages. The method is entirely inference-based: a single, unmodified SLM retrieves and follows instructions without additional training, and the approach generalizes across domains.
@@
We evaluate this framework across three reasoning benchmarks (MedQA, MMLU Law, and MathQA) to assess both its effectiveness and the conditions under which it provides the greatest benefit. Specifically, we investigate two questions. First, does instruction retrieval reliably improve the reasoning performance of small language models? Second, how does this benefit depend on model capacity and instruction design (e.g., length)? We show that across all three benchmarks, instruction retrieval improves accuracy by 5–10 percentage points for models with at least 3B parameters, without any additional fine-tuning Notably, on knowledge-intensive tasks such as MedQA and MMLU Law, a 14B-parameter model equipped with retrieved instructions surpasses GPT-4o in zero-shot accuracy. Further analysis shows that concise insteuctions yield the largest gains while performance remains stable even when instructions are shared across broader categories of problems. This robustness indicates that the approach does not depend on finely tailored instructions for each instance. To isolate the sources of improvement, we conduct a mixed-effects analysis and find that gains depend more on a model’s ability to follow structured guidance than on parameter count alone. In several cases, a smaller but more instruction-capable model outperforms a larger model with weaker underlying performance. These results suggest that externalizing reasoning as retrievable text offers a practical path to scaling reliable inference on resource-limited or privacy-sensitive hardware. All code and the Instruction Corpus will be released upon acceptance to support reproducibility.
@@ 2 Related Work
Chain of Thought. Chain-of-Thought (CoT) prompting encourages models to decompose a reasoning task into a sequence of intermediate steps rather than attempting to produce an answer directly. This approach has been shown to substantially improve the performance of LLMs across commonsense, symbolic, and mathematical reasoning benchmarks (Wei et al., 2023; Kojima et al., 2023; Wang et al., 2023). These gains, however, rely heavily on scale: only models with tens or hundreds of billions of parameters such as PaLM 540B (Chowdhery et al., 2022) or GPT-3 175B (Brown et al., 2020) , reliably benefit from CoT prompting. Smaller models frequently generate illogical reasoning traces, and their accuracy can even decline when forced to produce step-by-step rationales (Wei et al., 2023).
@@
High School Concise 453 4.64 4.83 …See the big clue (underweight athlete with missed periods), match it to the rule (FHA $\rightarrow$ low estrogen), then pick “low bone density”…
Graduate Concise 631 4.95 4.84 …Prioritize the diagnostic triad (amenorrhea + low BMI + training), apply HPO suppression logic, reject violated conditions (e.g., high TSH/estrogen), and select “decreased bone density”…
@@
Graduate Verbose 1,765 5.00 4.53 …Synthesize the amenorrhea–energy-deficit pattern, map it via leptin/kisspeptin $\rightarrow$ GnRH suppression to hypoestrogenism, weigh alternative axes (thyroid, prolactin, PCOS) by contradiction, and justify “decreased bone density” as the compelled outcome…
@@
Table 1: Comparison of four instruction variants differing in audience level and length. Columns report the average token length and Claude quality scores (five-point Likert) for knowledge comprehensiveness and relevance, with example excerpts from MedQA. Concise variants are 2–3× shorter while retaining high relevance, whereas verbose variants expand coverage at the cost of slightly lower relevance.
@@
Distillation. To overcome the scale dependence of prompting, a parallel line of work distills reasoning traces from LLMs into SLMs. These methods fine-tune smaller models on CoT-style rationales generated by larger models or derived from labeled data, with the goal of internalizing step-by-step reasoning skills (Ho et al., 2023; Li et al., 2022; Magister et al., 2023; Fu et al., 2023; Hsieh et al., 2023). Distillation has proven effective on arithmetic and symbolic reasoning benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), where reasoning patterns can be learned from repeated structures. However, the outcomes are less promising for knowledge-intensive tasks where accurate rationales depend on factual knowledge that smaller models struggle to retain. Because SLMs have limited parameter capacity, attempts to train them with extensive domain knowledge often lead to overspecialization and reduced generality (Fu et al., 2023).
@@
Retrieval. Rather than encoding all knowledge directly into model parameters, retrieval-based methods expand a model’s capacity by supplementing it with external information at inference time. Classical dense retrieval (Karpukhin et al., 2020) improves open-domain question answering by retrieving relevant passages, but the returned evidence is often noisy or unstructured, leaving the model to extract and organize reasoning steps on its own. Smaller models are more easily overloaded by long or noisy contexts and are less capable of abstracting structure from raw passages. As a result, the effectiveness of retrieval for SLMs depends on ensuring that the retrieved evidence is relevant and structured. More recent generate-then-retrieve frameworks address this limitation by synthesizing targeted context before retrieval (Yu et al., 2023; Wang et al., 2025), yielding more relevant evidence for reasoning. While these approaches have been developed for LLMs, their lessons are even more pertinent for SLMs.
@@
Knowledge Augmentation. Recent work (Kang et al., 2023; Zhao et al., 2024) improves on unstructured retrieval by more tightly coupling retrieved evidence with the reasoning process itself. Knowledge-Augmented Reasoning Distillation (Kang et al., 2023) grounds LLM-style rationales in retrieved passages and uses a reranker to emphasize evidence that directly supports those rationales. This helps smaller models learn reasoning that is explicitly tied to external knowledge, but it still relies on task-specific fine-tuning and encodes the resulting reasoning policy in model parameters. Probe then Retrieve and Reason (Zhao et al., 2024) modularizes this pipeline by separating it into two distilled SLMs: a probing model that identifies what knowledge is needed and formulates retrieval queries, and a reasoning model that constructs step-by-step rationales from the retrieved passages. While this separation improves interpretability, the reasoning procedure remains implicit in the model’s learned behavior and must be retrained to adapt to new domains.
@@ 3 Instruction Corpus
Small language models often fail on complex inference because they lack both sufficient domain knowledge and the structured support needed for multi-step reasoning. To address this gap, we build an Instruction Corpus, a collection of modular instructions for each domain. Each instruction has two parts: (i) background knowledge relevant to a problem type and (ii) step-by-step reasoning procedures for solving it. At inference, instructions are retrieved and included in the prompt, guiding reasoning without task-specific fine-tuning. For example, in MedQA, a group of training questions may ask about contraindications for anticoagulants. From this group, we derive an instruction that outlines how to check mechanisms of action, review contraindications, and compare relative risks.
Corpus Construction We construct the Instruction Corpus in three stages: clustering, instruction generation, and retrieval. First, training examples from each benchmark are embedded with OpenAI’s text-embedding-3-large and grouped using agglomerative clustering with average linkage and cosine distance. The resulting dendrogram supports different levels of granularity, where lower thresholds yield broad categories and higher thresholds produce highly specific clusters; by default, we adopt the most fine-grained partition so that each cluster corresponds to a distinct reasoning skill. Next, for each cluster, we generate a reusable instruction by prompting GPT-5 with standardized templates (D) and up to 5 question examples from the cluster. The input examples serve only to guide generation and do not appear verbatim in the final instructions. Finally, at inference time, test queries are embedded with the same encoder and matched by cosine similarity to the closest clusters; the top-$k$ instructions (default $k=5$) are retrieved and included in prompt to provide both factual grounding and procedural scaffolding for the given problem type. See Appendix 2 for threshold values and statistics.
@@
Instruction Variants While instructions provide missing knowledge and reasoning scaffolds, their effectiveness depends on how they are written. We therefore study how variation in audience level and length shapes small-model performance. For the audience level, we create two variants. A graduate-level version preserves domain terminology and assumes advanced conceptual familiarity (e.g., clinical reasoning in MedQA). The high-school-level version simplifies vocabulary and logical complexity (i.e., explain this to me like a high schooler). Aligning readability with prior knowledge reduces intrinsic cognitive load and improves comprehension, a core prediction of Cognitive Load Theory (Renkl and Atkinson, 2003; Sweller, 2024). Along the depth axis, we contrast concise instructions that contain only the essential reasoning steps with longer instructions that include additional commentary to stimulate germane load and schema formation. Moreover, recent prompt-engineering studies show that prompt compression often preserves, and sometimes even enhances, downstream accuracy despite large token reductions, which underscores the practical value of concise variants (Renze and Guven, 2024; Li et al., 2024). Crossing the two factors yields four prompt variants per task. Descriptive quality summaries by style are reported in Appendix Table 4. We also include a baseline prompt that does not enforce any constraints on style or length. This version reflects the instructions produced directly from our generation pipeline and serves as a neutral reference point. Comparing the four controlled variants against this baseline helps us to isolate the effects of readability and verbosity from artifacts of the generation process. Full prompt templates for all conditions, including the baseline, are provided in Appendix D.
@@
Quality Validation Not all instructions are equally useful. Generic prompts (e.g., ‘think carefully’) lack a stable, inspectable problem-solving procedure. By contrast, the documents in our corpus provide explicit domain background and concrete reasoning steps that define a consistent and inspectable decision process. Here, we verify that instructions are high quality and that the four variants differ only in style and length. To do so, we evaluate each instruction on three axes: knowledge (coverage and relevance of background facts), reasoning (soundness and task-specificity of steps), and clarity (structured, unambiguous presentation). A detailed rubric is given in Appendix A.1.3. Each document is scored on a five-point Likert scale by Claude Sonnet 4.5 to avoid same-model bias. We repeat evaluations three times and report averaged scores; full results are shown in Table 4. Across tasks, mean scores are consistently high ($\approx$4.6–5.0) on all three axes, confirming that the corpus provides accurate and reliable scaffolds rather than noisy or generic text. Table  1 summarizes quality differences across variants: verbosity increases comprehensiveness but slightly reduces relevance, while audience level has only minor effects. Clarity remains uniformly high across all conditions. Thus, the variants are equivalent in structure and quality, differing only in the controlled stylistic factors of length and audience.
@@
Corpus Profile The Instruction Corpus varies systematically in both length and size across tasks. Concise instructions average 500–750 tokens, while verbose instructions expand to 1.3k–2.1k tokens, a two- to threefold increase in prompt length. Length distributions are shown in Appendix Figure 6, with per-task breakdowns in Figure 7. Graduate-level instructions are slightly longer and more detailed than high-school versions, but this effect is small relative to the impact of verbosity. Corpus size is determined by the clustering granularity used during construction. At the default, most fine-grained threshold, the corpus contains approximately 29k instruction groups for MathQA, 8.4k for MedQA, and 1.0k for MMLU Law. Group sizes are highly skewed: most clusters contain only a single example, while the largest include up to 12. As a result, many instructions are narrowly scoped and grounded in one or a few examples, capturing the specific cues and reasoning steps needed for that problem type. Increasing the clustering threshold merges these singletons into larger groups, producing more general instructions that apply across broader classes of questions. We return to this tradeoff in Section 6, where we show that instruction retrieval remains robust across a wide range of cluster sizes. Full statistics and length distributions are reported in Appendix 2 and Appendix A.1.1.